### v1

# PEARL: Peer-Enhanced Adaptive Radio via On-Device LLM

Abstract
We present PEARL (Peer-Enhanced Adaptive Radio via On-Device LLM), a framework for cooperative cross-layer optimization in device-to-device (D2D) communication. Building on our previous work on single-device on-device LLMs (Lee et al., 2025), PEARL extends the paradigm by leveraging both publisher and subscriber states to guide Wi-Fi Aware (WA) parameter selection. A context-aware reward, which normalizes latency by application tolerances and modulates energy by device battery states, provides richer supervision for KL-based fine-tuning. We study two lightweight variants: PEARL (Head + Low-Rank Adaptation (LoRA) (Hu et al., 2022)) achieves the best overall performance, while PEARL-Lite (Head-only) delivers sub-20 ms inference at near-identical objective scores. Across synthetic scenarios grounded in real measurements, PEARL improves objective scores over heuristic and compact model baselines and reduces energy by up to 16% in cooperative low-battery cases. These results demonstrate that peer-aware context, reward-aligned training, and head-based efficiency make LLMs practical for always-on, on-device cross-layer control. Code, real-world demo, and dataset are available at https://github.com/abman23/pearl.
1 Introduction
Large language models (LLMs) are increasingly being adapted for resource-constrained edge devices (Qu et al., 2025), where they can enable adaptive networking, sensing, and cross-layer optimization (Shao et al., 2024) without relying on the cloud. Our previous work demonstrated that on-device LLMs can improve wireless performance by tuning cross-layer parameters in single-device settings, leveraging local context such as user location and time of day (Lee et al., 2025). While promising, this prior line of research leaves open the question of how LLM-based optimization can extend beyond isolated devices to cooperative, device-to-device (D2D) scenarios.
In D2D communication, link performance depends not only on the local device state but also on peer conditions such as battery level or device type. For example, when the publisher has ample energy but the subscriber is low on battery, an energy-preserving configuration may be preferable even at the cost of slightly higher latency. Designing a learning system that accounts for such asymmetric, two-sided context remains an open challenge. Moreover, efficient on-device deployment requires parameter-efficient fine-tuning (PEFT) strategies (Bucher and Martini, 2024; Han et al., 2024), but it is unclear how different PEFT modules compare in this setting, or how to ensure that training signals faithfully reflect heterogeneous application and device constraints.
Figure 1: System overview of PEARL. The agent runs on the publisher ($D_{1}$), combining local context (e.g., battery level, time, running applications) with peer-side context shared by the subscriber ($D_{2}$, e.g., battery level, device type). Based on these inputs, PEARL selects WA parameters $(\texttt{PerformanceMode},\texttt{AccessCategory})$, which configure the WA stack to minimize latency and energy on the D2D link.
Contributions. We propose PEARL (Peer-Enhanced Adaptive Radio via On-Device LLM), the first on-device LLM framework for cooperative cross-layer D2D optimization. Our key contributions are:
(i) Peer-context integration (Sec. 3.1, 4.3): subscriber-side information is incorporated into the model input, enabling two-sided adaptation beyond single-device optimization.
(ii) Context-aware reward (Sec. 3.2, 4.4): a reward function that normalizes latency by application-specific tolerances and modulates energy by device battery state, providing richer supervision for KL-based fine-tuning.
(iii) Head-based PEFT (Sec. 3.3, 4.6): lightweight classification heads (PEARL and PEARL-Lite) that achieve strong objective scores with sub-20 ms inference, narrowing the gap to the oracle while remaining deployable on mobile hardware.
Extensive evaluation on Wi-Fi Aware (Sec. 4 and Table 3) shows that PEARL consistently outperforms heuristic baselines and compact non-LLM models, demonstrating that peer-aware, reward-aligned training substantially improves both efficiency and robustness in D2D optimization.
2 Problem Formulation
Scenario. We consider a D2D link established via Wi-Fi Aware (WA) between a publisher $D_{1}$ and a subscriber $D_{2}$ (Camps-Mur et al., 2015). The publisher transmits application data while selecting cross-layer WA parameters that directly affect link performance. The decision model leverages both local and peer-side context, including device battery levels, device type (e.g., iPhone, iPad), time of day, and foreground application type sampled from a synthetic, time-varying distribution.
Action Space. At each decision step, the model chooses one of 8 discrete parameter tuples $(\texttt{PerformanceMode},\texttt{AccessCategory})$, where $\texttt{PerformanceMode}\in\{\texttt{realtime},\texttt{bulk}\}$ and $\texttt{AccessCategory}\in\{\texttt{bestEffort},\texttt{background},\texttt{interactiveVideo},\texttt{interactiveVoice}\}$.
Objective. The objective is to minimize link latency and device energy consumption while meeting application-specific requirements. Both local (publisher) and peer (subscriber) context guide parameter selection. For example, when the subscriber has low battery, the model should prefer energy-preserving modes even at the cost of slightly higher latency.
3 Method
3.1 System Architecture
PEARL builds on a pre-trained large language model (LLM) as a general-purpose feature encoder, and adapts it into a lightweight classifier for cross-layer parameter selection (Wu et al., 2024). Context features from both publisher and subscriber devices (e.g., battery levels, device type, time, application) are embedded into a structured prompt and passed to the frozen LLM backbone. A task-specific classification head is attached on top, predicting one of the discrete parameter tuples $(\texttt{PerformanceMode},\texttt{AccessCategory})$.
We consider two variants: PEARL (Head + Low-Rank Adaptation (LoRA) (Hu et al., 2022)), where the classification head is trained jointly with lightweight LoRA adapters inserted into the LLM, and PEARL-Lite (Head-only), where only the head is trained while the backbone remains frozen. For comparison, we also evaluate LoRA-only, where decisions are generated through the language modeling head without a classification head.
At inference, context features are encoded into a structured prompt, passed through the frozen LLM encoder, and mapped by the classification head to one of the 8 WA parameter tuples (Fig. 2).
Figure 2: Architecture of PEARL. Context features are encoded into a structured prompt and passed through a frozen LLM encoder. In PEARL-Lite, a classification head directly predicts one of 8 WA parameter tuples. In PEARL, LoRA adapters are added to the encoder and trained jointly with the head. In both cases, the head produces a single-token decision, avoiding autoregressive decoding.
3.2 Context-Aware Reward Design
We design a reward function that captures both application-specific latency requirements and device battery states. Each is normalized into a score, where higher values indicate better performance.
For an application $a$ with latency tolerance $L_{a}$ and parameters $p$ yielding average latency $\ell(p)$, the latency score is
$R_{\text{lat}}(a,p)=\max\!\left(100-100\cdot\tfrac{\ell(p)}{L_{a}},\,0\right).$
For a device $d$ with battery level $b_{d}$ and energy usage $E(p)$ under parameters $p$, the energy score is
$R_{\text{eng}}(d,p)=\frac{b_{d}}{E(p)}.$
The overall reward combines these terms across active applications $\mathcal{A}$ and devices $\mathcal{D}$:
$R(p)=w_{L}\cdot\tfrac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}R_{\text{lat}}(a,p)-w_{P}\cdot\tfrac{1}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\tfrac{1}{R_{\text{eng}}(d,p)}.$
Here, $\ell(p)$ and $E(p)$ are the measured latency and energy under parameters $p$, $\mathcal{A}$ is the set of active applications, and $\mathcal{D}=\{\text{publisher},\text{subscriber}\}$ the two devices considered. Weights $w_{L}$ and $w_{P}$ control the trade-off between latency and energy.
This formulation produces a reward distribution that yields soft labels for KL-based fine-tuning, preserving relative action preferences and improving generalization (Appendix C, H).
3.3 Post-Training Strategies
We evaluate several post-training strategies for adapting the LLM encoder and prediction head:
  • Supervised Fine-Tuning (SFT): We compare cross-entropy (CE) loss with Kullback–Leibler (KL) divergence loss. CE uses hard argmax labels, while KL leverages the full reward distribution as soft labels, preserving relative preferences across actions (Hinton et al., 2015).
  • Preference Optimization (DPO): Optionally, we apply Direct Preference Optimization (Rafailov et al., 2023) on paired contexts, where the higher-reward action is labeled as preferred and the lowest-reward action as non-preferred.
  • Head-based efficiency: In PEARL and PEARL-Lite, the classification head consumes the last hidden state and outputs logits over the 8 actions, producing a one-token decision. This avoids autoregressive decoding and yields sub-20 ms inference on edge hardware.
These strategies allow us to study the trade-offs between reward alignment, training cost, and inference efficiency (see Section 4).
4 Experiments
4.1 Setup
Dataset. We construct a simulation dataset from WA measurements combined with synthetic application usage distributions. For each parameter tuple $(\texttt{PerformanceMode},\texttt{AccessCategory})$, we measure average latency and power usage during controlled sessions. Each sample consists of $(\text{context},\text{action},\text{score})$, with scores computed as described in Section 3.2. Further details and parameter settings are provided in Table 4 and Appendix D.
Baselines. We evaluate our proposed framework for context-driven cross-layer optimization under the PEARL and PEARL-Lite variants introduced in Section 3.1. For comparison, we also include: (i) an oracle, which selects the global maximizer using ground-truth latency and energy; (ii) LoRA-only (LoRA); (iii) a rule-based baseline (Rule), which selects WA parameters heuristically based only on application type (see Table 8 in Appendix E); and (iv) two fixed baselines, Fixed (RT/IV) (fix-RT/IV, $(\texttt{realtime},\texttt{interactiveVoice})$) and Fixed (Bulk/BG) (fix-Bulk/BG, $(\texttt{bulk},\texttt{background})$).
Metrics. We report three metrics aligned with the definitions in Section 3.2: (i) the Objective Score, corresponding to the overall reward $R(p)$ that combines normalized latency and energy; (ii) Latency, the latency score $R_{\text{lat}}$ normalized by application-specific tolerance (higher is better); and (iii) Energy, the energy score $R_{\text{eng}}$, which reflects normalized energy consumption relative to device battery state (higher is better). Unless otherwise specified, results are averaged across scenarios covering different times of day and battery levels.
4.2 Performance Comparison
Figure 3: Objective score, latency, and energy comparison across PEARL variants and baselines.
Figure 3 shows that PEARL outperforms rule-based and fixed baselines on the joint objective score. The default PEARL variant achieves $7.58$, with PEARL-Lite close behind at $7.54$; both are about $7\%$ below the oracle ($8.16$) but remain well above LoRA-only ($7.32$) and Rule ($7.29$). While PEARL-Lite emphasizes lower latency at the cost of higher energy, PEARL strikes a more balanced trade-off, leading to the best overall performance. These results demonstrate that head-based designs enable effective context-driven adaptation and substantially narrow the gap to the oracle compared to heuristic strategies.
4.3 Impact of Peer Context Information
Variant Aggregated Scenario Cooperative Scenario (Low-Battery Subscriber)
Objective Score $\uparrow$ Latency $\uparrow$ Energy $\uparrow$ Objective Score $\uparrow$ Latency $\uparrow$ Energy $\uparrow$
w/ peer info (PEARL) 7.58 97.24 46.72 5.38 94.55 24.56
w/o peer info 7.55 97.98 44.50 5.04 99.13 20.53
w/ peer info (PEARL-Lite) 7.54 99.27 41.85 5.22 99.14 21.29
w/o peer info 7.49 99.27 40.99 5.03 99.13 20.47
Table 1: Impact of including peer information in the context. In the aggregated scenario, models trained w/ peer info achieve consistently higher objective scores than those w/o peer info. In the cooperative case (subscriber low battery), adding peer info reduces energy by up to $\sim$16% while maintaining latency, highlighting where peer awareness is most impactful.
Table 1 shows that incorporating peer (subscriber) information consistently improves performance. In aggregated scenarios, PEARL and PEARL-Lite w/ peer info achieve slightly higher objective scores than their w/o peer info counterparts. In the cooperative case, PEARL w/ peer info reduces energy consumption by $\sim$16% while maintaining latency within $0.5\%$ of w/o peer info, confirming that peer awareness provides tangible benefits in asymmetric conditions where coordination is critical. Additional qualitative snapshots illustrating this effect are provided in Appendix H.
4.4 Impact of Context-Aware Reward Design
Figure 4 shows that context-aware reward improves performance for both PEARL and PEARL-Lite. By normalizing latency against application-specific tolerances and incorporating device battery states into the energy term, the design generates richer supervision signals that better reflect heterogeneous requirements. Although absolute gains over the naive formulation are modest, they consistently yield higher objective scores, indicating more accurate adaptation. This demonstrates that careful reward design is crucial for leveraging context effectively in cross-layer optimization and for ensuring robustness across diverse usage scenarios.
Figure 4: Objective scores with context-aware vs. naive reward.
4.5 Effect of Post-Training Strategy
(a) PEARL
(b) PEARL-Lite
Figure 5: Objective scores for different training strategies. KL-based fine-tuning consistently outperforms CE and DPO, both in-distribution and out-of-distribution (OOD).
Figure 5 compares post-training strategies for PEARL and PEARL-Lite. Without fine-tuning (NoFT), objective scores are substantially lower in both in-distribution and OOD cases. Supervised fine-tuning with cross-entropy (SFT-CE) improves scores, while adding preference optimization (SFT+DPO) yields only marginal gains despite higher cost. In contrast, KL-based fine-tuning (SFT-KL) consistently achieves the best scores across both architectures. Its advantage stems from using soft labels derived from reward distributions, which preserve relative action preferences instead of collapsing them into hard labels. This provides richer supervision signals, enabling more nuanced trade-offs and better generalization across diverse scenarios.
4.6 Efficiency of Post-Training Modules
Baseline Objective Score $\uparrow$ Inference $\downarrow$ Training $\downarrow$
Time (ms) Tokens RAM (GB) RAM (GB) Time (min/epoch)
Base 7.04 760.7 (×1.0) 48.9 6.58
LoRA-only 7.32 243.1 (×3.1) 7.62 9.39 29.17 30
PEARL-Lite 7.54 14.5 (×52) 1.00 6.80 11.03 12
PEARL 7.58 40.3 (×18.9) 1.00 9.44 46.83 38
Table 2: Performance and efficiency of post-training modules. Parentheses denote speedup vs. Base (NoFT) . “Tokens” = average generated tokens per decision; head-based models always produce 1.00 (classification head). PEARL-Lite is most efficient, while PEARL yields the highest objective score.
Table 2 highlights the trade-off between performance and efficiency across different post-training modules. PEARL achieves the highest objective score ($7.58$) but incurs moderate training cost, while PEARL-Lite attains nearly comparable performance ($7.54$) with substantially lower inference latency ($14.5$ ms, $52\times$ faster than the Base) and reduced training memory. In contrast, LoRA-only yields lower objective scores ($7.32$) and slower inference, underscoring that head-based designs provide the most favorable balance for on-device deployment. This efficiency comes from the classification head: it consumes the last token’s hidden state to produce logits over the 8-action space (a single token), whereas LoRA-only relies on text generation, producing $\sim$8 tokens on average.
5 Discussion
Q1 — Interpretation: How do peer context and a context-aware reward improve PEARL’s on-device D2D decisions?. Incorporating subscriber-side context reduces partial observability and enables coordinated adaptation across devices. This leads to consistently higher objective scores, and in cooperative cases (e.g., low-battery subscriber) reduces energy by $\sim$16% with negligible latency change ($\leq$0.5%) (Table 1), showing that peer awareness makes PEARL more responsive to asymmetric conditions. The context-aware reward amplifies this effect by aligning training with heterogeneous constraints: latency is normalized by application-specific tolerances, and energy is scaled according to device battery states (Sec. 3.2, Fig. 4, Appendix C). Together, these choices guide PEARL toward energy-efficient yet latency-safe actions, improving robustness in asymmetric scenarios.
[…]
Q3 — Alternatives: Why not replace PEARL with a compact feed-forward network (FFN/MLP), a reinforcement-learning (RL) baseline, or a lightweight Transformer encoder?. While small feed-forward networks or RL baselines can be efficient for narrow tasks, they struggle with mixed-format context (numeric + categorical) and typically require additional feature engineering. Lightweight Transformer encoders such as MiniLM (22.7M parameters) (Wang et al., 2020) can process heterogeneous inputs more flexibly and achieve very low inference latency ($\sim$5–6 ms). However, as Table 3 shows, they yield lower objective and OOD scores than PEARL. By contrast, PEARL and PEARL-Lite leverage pretrained language modeling to naturally parse diverse context, providing stronger generalization while still running in the sub-20 ms range. Thus, although compact models may be attractive under extreme latency budgets, PEARL offers the best balance of accuracy, robustness, and deployability for always-on D2D optimization.
Baseline Objective Score $\uparrow$ OOD Score $\uparrow$ Inference Time (ms) $\downarrow$
PEARL 7.58 7.39 40.3
PEARL-Lite 7.54 7.39 14.5
MiniLM (Head+LoRA) 7.51 7.34 6.0
MiniLM (Head) 7.49 7.34 5.1
Table 3: Comparison of PEARL with a lightweight Transformer encoder (all-MiniLM-L6-v2, 22.7M).
Q4 — Limitations & Future Directions. We provide a proof-of-concept demo, with code, dataset, and a full demonstration video available at https://github.com/abman23/pearl to ensure reproducibility. Nevertheless, rigorous validation on hardware testbeds with real-time counters (e.g., per-packet latency and energy) is required for robust evaluation. Our study is currently limited to a narrow task space (2 PerformanceMode options $\times$ 4 AccessCategory options) and a single protocol (WA). Extending PEARL to richer parameter spaces, additional protocols (e.g., 5G), and heterogeneous devices remains an open challenge. Moreover, while head-based inference provides fast and accurate decisions, scaling to multiple tasks with separate heads may become burdensome. Future work should explore multi-task optimization strategies that preserve efficiency while maintaining accuracy.
Appendix A Real-World Demonstration
To validate the feasibility of PEARL beyond simulation, we implemented a real-world prototype as an iPadOS app. The app establishes D2D communication between two devices using Apple’s WA framework and integrates Apple’s on-device Foundation Models (AFM) to run the LLM agent directly on the publisher [Apple Inc., 2024]. The agent monitors local and peer-side context and dynamically updates WA parameters $(\texttt{PerformanceMode},\texttt{AccessCategory})$ to reconfigure the ongoing connection. We deploy the app on two iPad Pro devices (Apple M4 chip, 8 GB RAM, iPadOS 26.0). A full demonstration video is available at https://github.com/abman23/pearl.
Figure 6: Prototype demonstration of PEARL on two iPad Pro devices. The publisher (left) hosts the on-device LLM agent, while the subscriber (right) receives updated WA parameters.
Figure 7: Step-by-step illustration of WA parameter tuning. The publisher observes context features and proposes a new configuration $(\texttt{realtime},\texttt{interactiveVideo})$, then reconfigures the WA link and notifies the subscriber to update.
Appendix B Setup
Table 4 summarizes the hardware, software, dataset, and training configurations. For task design, we synthesize eight application types with time-of-day usage distributions that mimic typical mobile usage patterns (Fig. 8(a)). These distributions are used to sample application types as part of the context features in training and in-distribution evaluation. To test generalization, we construct an alternative distribution (Fig. 8(b)) to generate out-of-distribution (OOD) evaluation sets.
HW / SW
Device (Publisher / Subscriber) iPad Pro (M4, 8 GB RAM)
OS iPadOS 26.0
Frameworks Wi-Fi Aware, Foundation Models
LLM Backbone Llama-3.2-3B, AFM‑on‑device [Apple Inc., 2024]
Dataset (WA Log)
Total Samples 32,000
Train / Test Split 80% / 20%
Sampling Interval 5 s
Task Configuration
Action Space 2 modes × 4 categories = 8 actions
Scenarios 16 combinations of time-of-day (morning, afternoon, evening, night) and device battery levels (both high, medium, low, publisher high / subscriber low)
Application Types textMessage, voiceChat, videoCall, sensorSync, photoTransfer, videoUpload, firmwareUpdate, mapSync
Contextual Inputs Time, App Types, Publisher Battery Level, Subscriber Battery Level
Reward Weights $w_{L}=0.1$, $w_{P}=1.0$
History Window 10 past time steps
Training Configuration
Learning Rate $1\times 10^{-5}$
Batch Size / Grad. Accum. 4 / 16 (effective batch = 64)
Epochs 5
Weight Decay 0.01
Loss Functions CE, KL, DPO
LoRA Rank / Alpha 128 / 128
DPO Beta 0.1
Optimizer AdamW
GPU 1 × NVIDIA L40S
Table 4: Experiment setup including hardware, dataset, task configuration, and training details.
morning:
textMessage: 0.25,
voiceChat: 0.20,
videoCall: 0.25,
mapSync: 0.20,
photoTransfer: 0.10
afternoon:
textMessage: 0.20,
voiceChat: 0.20,
videoCall: 0.25,
sensorSync: 0.20,
photoTransfer: 0.15
evening:
textMessage: 0.10,
voiceChat: 0.20,
videoUpload: 0.30,
photoTransfer: 0.20,
videoCall: 0.20
night:
firmwareUpdate: 0.40,
sensorSync: 0.30,
textMessage: 0.10,
photoTransfer: 0.10,
mapSync: 0.10
(a) Training / in-distribution
morning:
videoCall: 0.30,
textMessage: 0.25,
voiceChat: 0.20,
photoTransfer: 0.15,
videoUpload: 0.10
afternoon:
videoCall: 0.25,
textMessage: 0.25,
voiceChat: 0.20,
photoTransfer: 0.20,
videoUpload: 0.10
evening:
videoUpload: 0.35,
videoCall: 0.25,
photoTransfer: 0.20,
voiceChat: 0.15,
textMessage: 0.05
night:
videoUpload: 0.30,
videoCall: 0.25,
textMessage: 0.20,
voiceChat: 0.15,
photoTransfer: 0.10
(b) Out-of-distribution evaluation
Figure 8: Synthetic application usage distributions used to generate context logs.
Appendix C Context-Aware Evaluation Metrics
We define two context-dependent metrics, Latency and Energy, to evaluate link quality under D2D dynamics. Unlike raw values, these metrics incorporate application- and device-specific constraints, providing not only quantitative measurement but also qualitative insight into how well system behavior aligns with user needs.
Latency is computed as a function of WA parameters and application type, expressed as a percentage score indicating how well the observed link latency satisfies the latency tolerance of the active application (Table 5). Energy is computed as a function of WA parameters and device battery level, reflecting how well the energy consumption of a configuration aligns with the current battery status.
Application Type Latency Tolerance (ms)
textMessage 200
voiceChat 50
videoCall 100
sensorSync 1,000
photoTransfer 2,000
videoUpload 5,000
firmwareUpdate 10,000
mapSync 500
Table 5: Latency tolerance thresholds for each application type.
To highlight the benefit of these metrics, we contrast them with raw measurements. Table 6 shows that while raw latency (ms) varies significantly across time periods, the Latency score remains consistent with application tolerance. For example, although average link latency at night (15.61 ms) is much higher than in the morning (4.97 ms), most nighttime applications are delay-tolerant, producing similarly high scores. Likewise, Table 7 compares Energy with raw energy consumption across high, medium, and low battery states. Here, the metric magnifies differences under low-battery conditions, assigning greater penalty and dynamically shifting the optimization objective toward energy efficiency.
Scenario (time, battery) Latency Score $\uparrow$ Link Latency (ms) $\downarrow$
(morning, both medium) 98.41 4.97
(afternoon, both medium) 98.43 6.60
(evening, both medium) 98.94 9.56
(night, both medium) 98.95 15.61
Table 6: Comparison of context-aware Latency score and raw link latency across scenarios.
Scenario (time, battery) Energy Score $\uparrow$ Energy Consumption (%/h) $\downarrow$
(afternoon, both high) 102.99 3.15
(afternoon, both medium) 59.46 3.24
(afternoon, both low) 24.26 2.90
Table 7: Comparison of context-aware Energy score and raw energy consumption across scenarios.
Appendix D Data Collection
The simulation dataset is constructed from real WA logs, combined with synthetic application usages defined in Fig. 8. To collect logs, we developed an iPadOS app that establishes a WA connection between two devices, simulates their interactions, and records context information. The app was deployed on two iPad Pro devices (Table 4), where we executed 16 sessions of data collection, each consisting of 2,000 consecutive logs. Each session corresponds to one specific scenario defined by time of day and device battery levels.
During each session, we iterated over all WA parameter tuples in the action space so that every configuration $(\texttt{PerformanceMode},\texttt{AccessCategory})$ was exercised. For each tuple, the app recorded both link latency values and battery usage, which are later used to compute the objective score as described in Section 3.2 and Section C.
Logged features. The following features were logged per record:
  • Device context: publisher and subscriber device IDs, battery levels, time of day, application type.
  • WA parameters: current PerformanceMode and AccessCategory.
  • Performance metrics: measured latency and battery usage.
Additional fields such as charging status, throughput capacity, signal strength, and throughput-related counters were also recorded (Figs. 9, 10), but were not used for training or inference. These remain available for future extensions or cross-validation.
Figure 9: Data collection pipeline. WA logs are recorded on the publisher and uploaded to a computer for pre-processing.
Figure 10: Example WA logs received by the computer, including device context, WA parameters, and performance metrics.
Appendix E Baseline Details (Rule)
The Rule baseline is a simple heuristic policy that depends only on application type. For each application, we assign a preferred WA parameter tuple based on common-sense intuition about its latency or throughput requirements (Table 8). At inference time, the baseline examines the past $n$ applications (with $n=10$ in our setup, consistent across all baselines) and selects the parameter tuple that appears most frequently among their preferred mappings.
This design makes Rule adaptive only to recent application history, without considering device battery states or other contextual factors. It thus serves as a lightweight non-learning baseline for comparison against context-aware methods such as PEARL and PEARL-Lite.
Application Type Preferred Parameter Tuple
textMessage (realtime, bestEffort)
voiceChat (realtime, interactiveVoice)
videoCall (realtime, interactiveVideo)
sensorSync (bulk, background)
photoTransfer (bulk, bestEffort)
videoUpload (bulk, bestEffort)
firmwareUpdate (bulk, background)
mapSync (realtime, bestEffort)
Table 8: Preferred WA parameter tuple assigned heuristically for each application type.
Appendix F Single-Objective Optimization Results
(a) Latency-only
(b) Energy-only
Figure 11: Performance comparison of PEARL variants and baselines for single-objective optimization.
Although PEARL is primarily designed for the joint optimization of latency and energy, we also evaluate performance under single-objective settings. Figure 11 compares PEARL variants against baselines when optimizing for Latency only or Energy only.
In the latency-only setting (Fig. 11(a)), PEARL, PEARL-Lite, and LoRA-only achieve scores within 1% of Fixed(RT/IV), which always selects $(\texttt{realtime},\texttt{interactiveVoice})$—a configuration strongly favoring low latency. In the energy-only setting (Fig. 11(b)), PEARL and PEARL-Lite achieve performance close to Fixed(Bulk/BG), the parameter choice that minimizes energy consumption, while outperforming other adaptive baselines.
Interestingly, PEARL-Lite performs on par with, and sometimes slightly better than, full PEARL in single-objective settings. This suggests that lightweight head-only models may be particularly well-suited for simpler optimization tasks, offering both efficiency and strong performance.
Appendix G Head Module Design Details
The head module in PEARL-Lite (and PEARL) is responsible for mapping the last-token embedding of the frozen LLM encoder to logits over the WA action space. Because this component is lightweight compared to the backbone LLM, its depth can be adjusted without incurring meaningful additional cost. We therefore explore different head architectures to examine whether deeper designs improve representational power.
Variant Performance $\uparrow$ Inference $\downarrow$ Training $\downarrow$
Objective Score Latency Energy Time (ms) RAM (GB) Time (min/epoch) RAM (GB)
3-layer Head 7.54 99.27 41.85 14.5 6.80 12 11.03
2-layer Head 7.50 99.27 41.21 14.6 6.80 12 10.54
1-layer Head 7.51 99.27 41.35 14.7 6.80 12 11.30
Table 9: Performance and efficiency of different head module depths in PEARL-Lite.
As shown in Table 9, the 3-layer head—the default configuration—achieves the highest objective score, while maintaining identical inference and training efficiency compared to shallower variants. This result suggests that slightly deeper heads can better utilize the LLM’s feature representations, while the overall system cost remains dominated by the backbone. Hence, deeper heads are a simple but effective design choice for improving performance without compromising efficiency.
Appendix H Full Scenario Snapshots
To complement the quantitative analysis in Section 4.3, we present two qualitative snapshots of context logs and the corresponding predictions made by PEARL, PEARL w/o peer info, and Rule. These examples illustrate how peer-side context guides parameter selection toward more energy-aware configurations when the subscriber is low on battery.
(a) Afternoon scenario.
(b) Evening scenario.
Figure 12: Context logs and decisions of PEARL, PEARL w/o peer info, and Rule under the subscriber low-battery case.
In Fig. 12(a), when the subscriber battery is only 10%, PEARL leverages this information to select (bulk, background), prioritizing energy saving. Without access to peer context, PEARL w/o peer info observes only the publisher’s high battery level and instead selects (realtime, bestEffort). Rule, based solely on application type, chooses (bulk, bestEffort), yielding the lowest objective score.
Fig. 12(b) shows a complementary case from the evening period. Here, despite a mix of latency-sensitive and high-throughput applications such as videoCall and videoUpload, PEARL again adapts to the low-battery subscriber and opts for (bulk, background). Both PEARL w/o peer info and Rule fail to adjust accordingly, resulting in lower scores.
Taken together, these snapshots demonstrate that PEARL consistently leverages peer-side context across different scenarios, leading to more energy-aware decisions while maintaining strong performance.

### v2

@@
 While promising, this prior line of research leaves open the question of how LLM-based optimization can extend beyond isolated devices to cooperative, device-to-device (D2D) scenarios.
-In D2D communication, link performance depends not only on the local device state but also on peer conditions such as battery level or device type.
+In D2D communication, link performance depends not only on the local device state but also on peer conditions such as battery level.
 For example, when the publisher has ample energy but the subscriber is low on battery, an energy-preserving configuration may be preferable even at the cost of slightly higher latency.
@@
 Objective.
-The objective is to minimize link latency and device energy consumption while meeting application-specific requirements.
+The objective is to balance link latency and device energy consumption while satisfying application-specific requirements.
 Both local (publisher) and peer (subscriber) context guide parameter selection.
@@
 PEARL builds on a pre-trained large language model (LLM) as a general-purpose feature encoder, and adapts it into a lightweight classifier for cross-layer parameter selection (Wu et al., 2024).
-Context features from both publisher and subscriber devices (e.g., battery levels, device type, time, application) are embedded into a structured prompt and passed to the frozen LLM backbone.
+Context features from both publisher and subscriber devices (e.g., battery levels, time, application) are embedded into a structured prompt and passed to the frozen LLM backbone.
 A task-specific classification head is attached on top, predicting one of the discrete parameter tuples $(\texttt{PerformanceMode},\texttt{AccessCategory})$.
@@
 We evaluate our proposed framework for context-driven cross-layer optimization under the PEARL and PEARL-Lite variants introduced in Section 3.1.
-For comparison, we also include: (i) an oracle, which selects the global maximizer using ground-truth latency and energy; (ii) LoRA-only (LoRA); (iii) a rule-based baseline (Rule), which selects WA parameters heuristically based only on application type (see Table 8 in Appendix E); and (iv) two fixed baselines, Fixed (RT/IV) (fix-RT/IV, $(\texttt{realtime},\texttt{interactiveVoice})$) and Fixed (Bulk/BG) (fix-Bulk/BG, $(\texttt{bulk},\texttt{background})$).
+For comparison, we also include: (i) an oracle, which selects the configuration that maximizes the context-aware reward derived from ground-truth latency and energy — effectively representing the optimal balance between latency and energy consumption under the defined application constraints; (ii) LoRA-only (LoRA); (iii) a rule-based baseline (Rule), which selects WA parameters heuristically based only on application type (see Table 8 in Appendix E); and (iv) two fixed baselines, Fixed (RT/IV) (fix-RT/IV, $(\texttt{realtime},\texttt{interactiveVoice})$) and Fixed (Bulk/BG) (fix-Bulk/BG, $(\texttt{bulk},\texttt{background})$).
 Metrics.
@@
 In the aggregated scenario, models trained w/ peer info achieve consistently higher objective scores than those w/o peer info.
-In the cooperative case (subscriber low battery), adding peer info reduces energy by up to $\sim$16% while maintaining latency, highlighting where peer awareness is most impactful.
+In the cooperative scenario—where the subscriber is low on battery and the publisher adapts its configuration to assist—adding peer info reduces energy consumption by up to $\sim$16% while maintaining reasonable latency, highlighting where peer awareness provides the greatest benefit.
 Table 1 shows that incorporating peer (subscriber) information consistently improves performance.
@@
 Each session corresponds to one specific scenario defined by time of day and device battery levels.
+Battery levels were sampled uniformly across high, medium, and low states to ensure balanced representation of energy conditions.
 During each session, we iterated over all WA parameter tuples in the action space so that every configuration $(\texttt{PerformanceMode},\texttt{AccessCategory})$ was exercised.
@@
 It thus serves as a lightweight non-learning baseline for comparison against context-aware methods such as PEARL and PEARL-Lite.
+Future extensions could incorporate a simple “battery-safe” heuristic—e.g., switching to energy-preserving modes below 20% battery—to better reflect real-world smartphone behavior.
 Application Type Preferred Parameter Tuple